A supply plane, flying 1.5 km above the ground drops food supply to a certain barangay below. If a bag of

food supply is 5 kg , find

a) the potential energy of a food supply bag just before it was drop ?

b) the potential energy of the supply bag after 8 sec.

c) calculate work done by the gravity after 8 sec


1
Expert's answer
2021-05-06T13:16:11-0400

a)


"U=mgh=(5)(9.8)(1500)=73500\ J"

b)


"U'=mg(h-0.5gt^2)\\=(5)(9.8)(1500-0.5(9.8)(8)^2)=58134\ J"

c)


"W=58134-73500=-15366\ J"